a certain mountain has has an elevation of 19260. in 1916 the glacier on this peak covers 9 acres.by 2001 the glacier has melted to only 1 acre.
what is the equation that gives the acreage of the glacier t years after 1916?

Answer by stanbon(75887)

a certain mountain has has an elevation of 19260. in 1916 the glacier on this peak covers 9 acres.by 2001 the glacier has melted to only 1 acre.
what is the equation that gives the acreage of the glacier t years after 1916?
---------------------------------------------
Let "t" be the numbre of years after 1916.
You have 2 points: (0,9) and (85,1)
----
slope: m = (1-9)/(85-0) = -8/85 = -0.094
intercept: b = 9
----------------
Equation:
area = -0.094t+9
=================
Cheers,
Stan H.
=====================